# A semi-analytical solution for the lubrication force between two spheres approaching in viscoelastic fluids described by the Oldroyd-B model under small Deborah numbers

### v1

@@ 2.4 Dimensionless equations
The system of partial differential equations to solve is given by (6)-(8), and (11)-(14). Under the assumption of very small interparticle gaps, one can assume that the relation between the gap and the reduced radio is small, i.e., $\epsilon=\frac{h_{0}}{a}\ll 1$; this allows to adopt an order-of-magnitude analysis of the system of partial differential equations through lubrication theory (Kim and Karrila, 1991). With this aim, we start by non-dimensionalising the system above in non-dimensional stretched coordinates $\{R,Z\}$ (Jeffrey, 1982; Rodin, 1996):
@@
where the approaching velocity $V$ is the characteristic velocity in the $z-$direction, and $U=\frac{V}{\epsilon^{1/2}}$ is the characteristic velocity in the $r-$direction (direction of the flow process), derived from (6). Here, the dimensionless group-numbers of Deborah ($De$) and Reynolds ($Re$) measure the ratio between elastic and viscous forces, and the ratio between inertial and viscous forces, respectively. Moreover, since the Re number between particles is small, temporal and convective terms may be ignored in the momentum balance. Hence, the conservation equations (6)-(8) appear in dimensionless form, as follows:
@@
To show explicitly the effects of the viscoelasticity through normal-stress differences that may arise from viscoelastic-stress, a re-casted expression for the momentum balance equation in r-direction can be obtained by adding and subtracting $\frac{S_{zz}}{R}$ on (17), which then may be rearranged as:
@@
$S_{rz}+De\left(\frac{\partial S_{rz}}{\partial T}+u\frac{\partial S_{rz}}{\partial R}+w\frac{\partial S_{rz}}{\partial Z}-S_{rz}\left(\frac{\partial u}{\partial R}+\frac{\partial w}{\partial Z}\right)-\frac{S_{zz}}{\epsilon^{1/2}}\frac{\partial u}{\partial Z}-\right.\\ \left.\epsilon^{1/2}S_{rr}\frac{\partial w}{\partial R}\right)=(1-\beta)\left(\epsilon\frac{\partial w}{\partial R}+\frac{\partial u}{\partial Z}\right),$ (21)
@@ 3 Steady-state case
Under the assumption of spheres being very close between each other, i.e., $\epsilon=\frac{h_{0}}{a}\ll 1$, the terms in (20)-(23) with order-of-magnitude $O(\epsilon)^{1/2}$ and higher may be considered negligible. In this first section, we also consider a steady-state approximation, making null the temporal partial-derivative terms, i.e., $\frac{\partial(\cdot)}{\partial T}=0$ in the viscoelastic-stress. Moreover, to get a first approximation, we consider a weakly-viscoelastic instance through a small De-numbers limiting case, considering De-numbers with a similar magnitude to $\epsilon^{1/2}$, such that $\frac{De}{\epsilon^{1/2}}\sim O(1)$. Under such set of assumptions, the viscoelastic-stress components reduce to:
@@
and, therefore, $N_{1}=S_{rr}$ and $N_{2}=0$. The first normal-stress difference is led by the normal stress in the flow direction, i.e., $S_{rr}$. Conversely, the second normal-stress difference is zero and it will have no influence in the motion of the fluid. Then, the momentum balance equations are obtained by substituting (28) and (29) in (18) and (19), such that:
@@
From examination of (30) and (31), one may conclude that the pressure $P$-dependence appears only in the radial direction. In (30), the viscous response is given by the shear-stress, such that $\frac{\partial^{2}u}{\partial Z^{2}}$ manifests the viscous resistance to the rate-of-deformation. Complementarily, the viscoelastic response is led by the first normal-stress difference, such that $\frac{2(1-\beta)De}{R}\left(\frac{\partial u}{\partial Z}\right)^{2}$, represents a measure of the resistance of deformation. This viscoelastic term in (30) represents an additional resistance to shearing deformation that the fluid exhibits in the radial direction, that results from the stretching and alignment of the material along the streamlines (Bird et al., 1987), and which radial pressure gradient must overcome.
@@
where, $\alpha=\frac{2(1-\beta)De}{R}$ and $\omega=\sqrt{-\alpha\frac{\partial P}{\partial R}}$, and whose solution step-by-step is provided in Appendix A. Analytic solution for velocity profile. Here, the pressure gradient is assumed to be negative and is yet to be determined. The velocity profile in expression (32) draws a parabolic profile at very low De-numbers; however, due to the first normal stress differences ($N_{1}$) and geometric changes trough $\left(\frac{N_{1}}{R}\right)$ in (30) the velocity profile is reduced in the centre of the gap between the spheres. Figure 2 shows the influence of viscoelasticity in the radial velocity profile under steady-state. We introduce a normalised velocity defined as $\frac{u}{u_{Newt}}$ that represents the relation of the viscoelastic velocity profile with respect to Newtonian velocity profile under the same pressure gradient. The Newtonian velocity profile is recovered from solution (32) when taking $De=0$ and $\beta=1$, such that:
@@
A viscoelastic solution characterized by $De=10^{-6}$ behaves as (close to) Newtonian. By increasing De-number to $\{10^{-3},10^{-2}\}$ the solutions display negligible variation with respect to the Newtonian case along the axial direction. However, for the $De=10^{-1}$ case, a slight difference is apparent around the maximum velocity (see inset in figure 2), which is situated below the Newtonian profile. This result suggests that viscoelasticity slows the velocity distribution at the centreline relative to the Newtonian profile under the same pressure gradient.
Figure 2: Normalised radial velocity profile for a Newtonian fluid and viscoelactic fluid characterized by the Oldroyd-B model with $De=\{10^{-3},10^{-2},10^{-1}\}$ and $\beta=0$.
@@ 3.1 Flow rate and mass conservation
Flow rate, $q$, arises when two approaching spheres squeeze the fluid trapped between them, and is given by the relative movement of the particles, the dimensionless flow rate, $Q$, renders as follows:
@@
The expression (34) indicates that the flow-rate is a function of the radial direction, therefore, the driving-force that drives the velocity profile (pressure gradient in this case) must have a determined value in order to satisfy the mass conservation. Integrating the mass conservation equation (Hori, 2006; Ruangkriengsin et al., 2024) in (16) under the boundary condition in (9), in proper dimensionless form, i.e. $w(Z_{1})=0$ and $w(Z_{2})=-1$, it can be written as:
@@
Then, (36) establishes the mass-conservation condition that will be employed for subsequent calculations (see Appendix B. Continuity equation for further details). The steps to determine velocity profile, pressure gradient and solution of (36) at each radial position, from the centreline ($R_{i}=0$) to the edge of the spheres ($R_{i}=a$), assuming spheres of the same radius $\left(\kappa_{a}=\frac{1}{2}\right)$, are given as follows:
  • First, for a given radial position ($R_{i}$), two estimates of $\frac{\partial P}{\partial R}_{i}$ are proposed from the Newtonian solution: $\frac{\partial P}{\partial R}_{New}=\frac{-6R}{H^{3}(R)}$, which is equal to the expression given in Vázquez-Quesada and Ellero (2016) in dimensionless form.
  • Secondly, $Q$ is calculated for both estimates. Here, the integral on the right-hand-side of (34) is determined by applying the Simpson’s 3/8 rule.
@@
The pressure gradient profile in the radial direction is illustrated in figure 3(a). The pressure gradient shows a maximum near the centreline, beyond which it drops steeply along radial direction. The pressure gradient increases slightly for $De=\{10^{-3},10^{-2}\}$; however, for $De=10^{-1}$ it can be up to $\approx 4\%$ higher than in the Newtonian case at the peak.  Then, the pressure can be approximated by integrating the pressure gradient from the figure 3(a) within integration limits ranging from $R=0$ to $R=a$ using the Euler’s method, to calculate the expression:
@@
Pressure attains its maximum in the centreline, located at $R=0$, and declines rapidly within a small radius range, as illustrated in figure 3(b). Solutions under small De-numbers $De=\{10^{-3},10^{-2}\}$ seem similar to the Newtonian-solution. Only under $De=10^{-1}$, there is an evident deviation from the Newtonian behaviour, $\approx 6\%$ higher than in the Newtonian case, at the peak. These findings may be explained from an analysis of (30) and (36), where one may find an additional force from the viscoelastic characteristics of the fluid, resulting from the first normal stress-difference. This is indicated by a slight reduction in velocity profile (figure 2), which consequently leads to a decrease in flow rate as described in (34). Therefore, to maintain mass conservation detailed in (36), there is a corresponding rise in both the pressure gradient and pressure.
Figure 3: Numerical solutions for (a) the pressure gradient (Simpson’s 3/8 rule and secant method) and (b) pressure (Euler’s method) profiles under the viscoelastic parametrisation: $\beta=0$, $a=1$, $h_{0}=2.5\times 10^{-3}$ and $De=\{10^{-3},10^{-2},10^{-1}\}$ under the same approach velocity of the sphere.
@@ 3.2 Stress and normal forces
The normal force acting on the particles’ surface, $f$, is therefore obtained by integration of the total stresses along their surface. Again, Euler’s method is employed to calculate the force from the numerical approximation of pressure in the range of $R=0$ to $R=a$. The dimensionless net viscoelastic lubricating force normal, $F$, to the approach velocity direction on the particle surface $S_{2}$ is given by:
@@
Figure 4 compares the percentage increase of viscoelastic lubrication force to that of Newtonian lubrication force, given in dimensionless form such as: $F_{Newt}=\frac{1}{4\epsilon}$. As illustrated in figure 4, the lubrication force exerted on the particle surface $S_{2}$ increases weakly due to increases in the pressure gradient caused by the viscoelastic response. Notably, the increasing presence of the Newtonian solvent in the dissolution, measured under a $\beta$-increase, reduces the force intensity. For $De=1\times 10^{-1}$ and $\beta=0$, the viscoelastic lubrication force shows the greatest rise in relation to the Newtonian lubrication force, reaching an increase of up to $\approx 2\%$. In a steady state, the lubrication force exerted on the particles is immediate. As discussed in this section, the incorporation of viscoelasticity into the momentum balance equation (via $N_{1}$) results in minor effects in the velocity, pressure gradient, and pressure distribution. Thus, the effect of viscoelasticity on the lubrication force is small when the fluid response is considered instantaneous and when the effects of geometric changes derived from the particle approach are not considered.
Figure 4: Percentage increase on the viscoelastic force with respect to a Newtonian response under different $\beta$ for different De-numbers.
@@ 4.1 Transient analysis
The steady-state case demonstrates the instantaneous response of viscoelastic fluids when subjected to a constant squeezing velocity. Nevertheless, because of particle dynamics and the time response of the viscoelastic fluids, we introduce a time-dependent scenario. Therefore, we analyse the transient case by retaining the time-dependent terms only in the stress components (11)-(14), which renders the following partial differential-equation system:
@@
We consider a completely relaxed stress for the initial conditions, as follows:
@@
As the particle $S_{2}$ moves with constant velocity $V$ towards the particle $S_{1}$, its contour change with time, such as: $z_{2}(r,t)=h_{0}+\frac{r^{2}}{2a}-Vt$. Therefore, the interparticle distance in dimensionless form reduces with time as following:
$H(R,T)=\frac{z_{2}(r,t)-z_{1}(r)}{h_{0}}=\frac{h(r,t)}{h_{0}}=1+\frac{R^{2}}{2\kappa_{a}}-T,$ (44)
@@ 4.2 Transient solutions
Substituting (46) in (40) and solving for the dimensionless shear stress under initial condition in (43), one gathers:
@@
Differentiation with respect to $Z$ of (49) and grouping similar terms, and considering:
@@
momentum balance equation in $r-$direction results in:
@@
Momentum equation in (51) relies on temporal functions $\left(\alpha(T),\frac{\partial P^{\prime}}{\partial R}(T)\right)$ that come from the viscoelastic-stress. These terms emerge from the time-dependent development of the velocity field, as $S_{2}$ contour reshapes over time. Integration of (51) under the boundary condition in (9), yields the following velocity profile:
@@
Here, the temporal functions have the following time-limiting values:
@@
From the temporal functions limits in (53) and (54), we may conclude that:
@@
    which comes from (51) when viscoelastic effects are zero and represents the force balance for a Newtonian fluid of viscosity $\eta_{0}=\eta_{s}$. Therefore, the viscoelastic behaviour starts from a Newtonian fluid response.
@@
  • In the range $10<T<$ 1, the viscoelastic fluid response may evolve from a Newtonian fluid of viscosity $\eta_{0}=\eta_{s}$ to a viscoelastic fluid of viscosity $\eta_{0}=\eta_{s}+\eta_{p}$, whose force balance is given by (51). The dimensionless time expected to reach a steady value of $(g(T),f(T))$ is around $T\sim 0.5$ even for $De=1\times 10^{-1}$ (see figure 5). However, since $Z_{2}(R,T)$ varies with time, the distance between the spheres will diminish and generate geometric changes (Ruangkriengsin et al., 2024), indicating that they are moving closer to touching each other.
@@
  • In the limit of $T\rightarrow 1$, (54) simplifies to a version resembling the steady-state scenario in (30), that is, a viscoelastic fluid with a viscosity of $\eta_{0}=\eta_{s}+\eta_{p}$. Nevertheless, the contour of sphere $S_{2}$ continues to draw nearer to sphere $S_{1}$, which promotes the fluid to get confined.
@@
Figure 5: Temporal function from viscoelastic-stress for $De=\{1\times 10^{-2},1\times 10^{-1}\}$ and $\beta=0.5$.
@@
Therefore, the response at relatively short times resembles the response of a Newtonian fluid, indicating that the parameter $\beta$ modulates the initial lubrication force. In the case of an upper-convected Maxwell fluid, i.e., for $\beta=0$, the force will start from zero, and, for $\beta=1$, the force will align with the expected reaction of a Newtonian fluid.  The computations for pressure gradient, pressure, and lubrication force follow the same numerical techniques outlined in Section 3.1 adding the functionality with time through the terms $Z_{2}(R,T)$, $g(T)$ and $f(T)$. Despite the steady-state problem, given the time in the domain $(0\leq T<1)$, the viscoelastic lubrication force is expected to increase over time or increase as distance is reduced, according to (44). In order to detect a significant alteration in distance, time needs to exceed $0.9$, such as $T>0.9$.
@@
We compare our solutions with expressions (52) and (54) in Zheng et al. (2023), corresponding to the pressure and the lubrication force, respectively. The principal distinction between our formulation and that of Zheng et al. (2023) lies in their use of a Newtonian velocity profile and a stress-based boundary condition to compute the pressure gradient. Figure 6 presents the time evolution of the pressure for different values of De-numbers. For early times, specially when $T\lesssim 0.3$, our predictions yields pressures lower than those of the Newtonian reference solution. However, as $T\gtrsim 0.3$, the viscoelastic predictions converge to, and marginally exceed, the Newtonian result. In that sense, the development of the viscoelastic time-dependent functions in (51) rise both pressure and lubrication force. A notable divergence from (52) becomes apparent as $T\rightarrow 1$. Their model predicts a decline in pressure, approaching negative values, in contrast to the monotonic increase characteristic of the Newtonian solution.  For the lubrication force in figure 7, our solutions are in close agreement with the analytical predictions of Zheng et al. (2023) at short times. Also, both predictions continue in close agreement with the Newtonian upper limit for all selected values of De-numbers. Nonetheless, for $T\gtrsim 0.2$, Zheng’s predictions indicate a lubrication force that exceeds our predicted values.
@@
Figure 6: Dimensionless pressure against time at different values of $De=\{1\times 10^{-6},5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},1\times 10^{-1}\}$ for $\epsilon=2.5\times 10^{-3}$. Symbols are our numerical solutions and lines come from (52) in Zheng et al. (2023).
@@
Figure 7: Dimensionless force against time at different values of $De=\{1\times 10^{-6},5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},1\times 10^{-1}\}$ for $\epsilon=2.5\times 10^{-3}$. Symbols are our numerical solutions and lines come from (52) in Zheng et al. (2023).

### v2

@@ 2.4 Dimensionless equations
The system of partial differential equations to solve is given by (6)-(8), and (11)-(14). Under the assumption of very small interparticle gaps, one can assume that the ratio between the gap and the reduced radius is small, i.e., $\epsilon=\frac{h_{0}}{a}\ll 1$; this allows to adopt an order-of-magnitude analysis of the system of partial differential equations through lubrication theory (Kim and Karrila, 1991). With this aim, we start by non-dimensionalising the system above in non-dimensional stretched coordinates $\{R,Z\}$ (Jeffrey, 1982; Rodin, 1996):
@@
where the approaching velocity $V$ is the characteristic velocity in the $z-$direction, and $U=\frac{V}{\epsilon^{1/2}}$ is the characteristic velocity in the $r-$direction (the direction of the flow process), derived from (6). Here, the dimensionless group-numbers of Deborah ($De$) and Reynolds ($Re$) measure the ratio between elastic and viscous forces, and the ratio between inertial and viscous forces, respectively. Moreover, since the Re number between particles is small, temporal and convective terms may be ignored in the momentum balance. Hence, the conservation equations (6)-(8) appear in dimensionless form, as follows:
@@
To show explicitly the effects of the viscoelasticity through normal-stress differences that may arise, a re-casted expression for the momentum balance equation in the r-direction can be obtained by adding and subtracting $\frac{S_{zz}}{R}$ on (17), which then may be rearranged as:
@@
$S_{rz}+De\left(\frac{\partial S_{rz}}{\partial T}+u\frac{\partial S_{rz}}{\partial R}+w\frac{\partial S_{rz}}{\partial Z}-S_{rz}\left(\frac{\partial u}{\partial R}+\frac{\partial w}{\partial Z}\right)-\frac{S_{zz}}{\epsilon^{1/2}}\frac{\partial u}{\partial Z}-\epsilon^{1/2}S_{rr}\frac{\partial w}{\partial R}\right)=\\ (1-\beta)\left(\epsilon\frac{\partial w}{\partial R}+\frac{\partial u}{\partial Z}\right),$ (21)
@@ 3 Steady-state case
Under the assumption of spheres being very close between each other, i.e., $\epsilon=\frac{h_{0}}{a}\ll 1$, the terms in (20)-(23) with order-of-magnitude $O(\epsilon)^{1/2}$ and higher may be considered negligible. In this first section, we also consider a steady-state approximation, making null the temporal partial-derivative terms, i.e., $\frac{\partial(\cdot)}{\partial T}=0$ in the viscoelastic stress. Moreover, to get a first approximation, we consider a weakly-viscoelastic instance through a small De-number limiting case, considering De-numbers with a similar magnitude to $\epsilon^{1/2}$, such that $\frac{De}{\epsilon^{1/2}}\sim O(1)$. Under such set of assumptions, the viscoelastic-stress components reduce to:
@@
and, therefore, $N_{1}=S_{rr}$ and $N_{2}=0$. The first normal-stress difference is led by the normal stress in the flow direction, i.e., $S_{rr}$. Conversely, the second normal-stress difference is zero and it will have no influence in the motion of the fluid. Then, the momentum balance equations are obtained by substituting (28) and (29) in (18) and (19), respectively, such that:
@@
From examination of (30) and (31), one may conclude that the pressure $P$-dependence appears only in the radial direction. In (30), the viscous response is given by the shear-stress, such that $\frac{\partial^{2}u}{\partial Z^{2}}$ manifests the viscous resistance to the rate-of-deformation. Complementarily, the viscoelastic response is led by the first normal-stress difference, such that $\frac{2(1-\beta)De}{R}\left(\frac{\partial u}{\partial Z}\right)^{2}$, represents a measure of the viscoelastic resistance of deformation. This viscoelastic term in (30) represents an additional resistance to shearing deformation that the fluid exhibits in the radial direction, that results from the stretching and alignment of the material along the streamlines (Bird et al., 1987), and which the radial pressure gradient must overcome.
@@
where $\alpha=\frac{2(1-\beta)De}{R}$ and $\omega=\sqrt{-\alpha\frac{\partial P}{\partial R}}$, and whose solution step-by-step is provided in Appendix A. Here, the pressure-gradient is assumed to be negative and is yet to be determined. The velocity profile in expression (32) draws a parabolic functionality at very low De-numbers; however, due to the first normal-stress difference ($N_{1}$) and geometric changes through $\left(\frac{N_{1}}{R}\right)$ in (30), the velocity profile is reduced in the centre of the gap between the spheres. Figure 2 shows the influence of viscoelasticity in the radial velocity profile under steady-state. We introduce a normalised velocity, defined as $\frac{u}{u_{Newt}}$ that represents the ratio of the viscoelastic velocity profile with respect to Newtonian velocity profile under the same pressure-gradient. The Newtonian velocity profile is recovered from solution (32) when taking $De=0$ and $\beta=1$, such that:
@@
A viscoelastic solution characterised by $De=10^{-6}$ behaves close to the Newtonian case. By increasing viscoelasticity to $De=\{10^{-3},10^{-2}\}$, solutions display negligible variation with respect to the Newtonian case along the axial direction. However, for the $De=10^{-1}$ case, a slight difference is apparent around the maximum velocity (see inset in Figure 2), which lies below the Newtonian profile. This result suggests that viscoelasticity slows down the velocity at the centreline for the viscoelastic case, this being relative to the Newtonian profile under the same pressure-gradient.
Figure 2: Normalised radial velocity profile for a Newtonian fluid and a viscoelastic fluid characterised by the Oldroyd-B model under $De=\{10^{-3},10^{-2},10^{-1}\}$ and $\beta=0$.
@@ 3.1 Flow rate and mass conservation
The flow rate $q$ arises when two approaching spheres squeeze the fluid trapped between them, and is provoked by the relative motion of the particles. The dimensionless flow-rate measure $Q$ is defined as follows:
@@
The expression (34) indicates that the flow-rate is a function of the radial direction; therefore, the force driving the velocity profile, i.e., the pressure-gradient in this case, must have a determined value in order to satisfy the mass conservation principle. Integrating the mass conservation equation (Hori, 2006; Ruangkriengsin et al., 2024) in (16) under the boundary condition in (9), in proper dimensionless form, i.e. $w(Z_{1})=0$ and $w(Z_{2})=-1$, yields:
@@
Then, (36) establishes the mass-conservation condition that will be employed for subsequent calculations (see Appendix B for further details). The steps to determine the velocity profile, the pressure-gradient and the solution of (36) at each radial position, from the centreline ($R_{i}=0$) to the edge of the spheres ($R_{i}=a$), assuming spheres of the same radius $\left(\kappa_{a}=\frac{1}{2}\right)$, are given as follows:
  • Firstly, for a given radial position ($R_{i}$), two estimates of $\frac{\partial P}{\partial R}_{i}$ are proposed from the Newtonian solution, i.e., $\frac{\partial P}{\partial R}_{New}=\frac{-6R}{H^{3}(R)}$, which is equivalent to the expression given in Vázquez-Quesada and Ellero (2016) in dimensionless form.
  • Secondly, $Q$ is calculated for both estimates. Here, the integral on the LHS of (34) is determined by applying the Simpson’s 3/8 rule.
@@
The pressure-gradient profile in the radial direction is illustrated in Figure 3(a). The pressure gradient shows a maximum near the centreline, beyond which it drops steeply along the radial direction. The pressure-gradient increases slightly for $De=\{10^{-3},10^{-2}\}$; however, for $De=10^{-1}$, it can be up to $\approx 4\%$ higher than in the Newtonian case at the peak.
Then, the pressure can be approximated by integrating the pressure-gradient from Figure 3(a) within integration limits ranging from $R=0$ to $R=a$, using Euler’s method, to obtain:
@@
Pressure attains its maximum in the centreline, located at $R=0$, and declines rapidly within a small radius range, as illustrated in Figure 3(b). Solutions under small De-numbers $De=\{10^{-3},10^{-2}\}$ appear similar to the Newtonian solution. Only under $De=10^{-1}$, there is an evident deviation from the Newtonian behaviour, with pressures some $\approx 6\%$ higher than in the Newtonian case at the peak. These findings may be explained from an analysis of (30) and (36), where one finds an additional force originating from viscoelasticity and the first normal-stress difference. This is correlated with a slight reduction in the velocity profile (Figure 2), which consequently leads to a decrease in flow rate as described in (34). Therefore, to comply with mass conservation (36), there is a corresponding rise in both pressure gradient and pressure.
Figure 3: Numerical solutions for: (a) the pressure-gradient (Simpson’s 3/8 rule and secant method) and (b) pressure (Euler method) under the viscoelastic parametrisation: $\beta=0$, $a=1$, $h_{0}=2.5\times 10^{-3}$, $De=\{10^{-3},10^{-2},10^{-1}\}$, and under the same sphere approach velocity.
@@ 3.2 Stress and normal forces
The normal force acting on the particles’ surface ($f)$ is therefore obtained by integration of the total stresses along their surface. Again, Euler’s method is employed to calculate the force from the numerical approximation of pressure in the range of $R=0$ to $R=a$. The dimensionless net viscoelastic lubricating force ($F$) normal to the approach-velocity direction on the particle surface $S_{2}$ is given by:
@@
Figure 4 compares the percent increase of the viscoelastic lubrication force against that of Newtonian lubrication force; such Newtonian measure is defined in dimensionless form as: $F_{Newt}=\frac{1}{4\epsilon}$. As illustrated in Figure 4, the normal lubrication force exerted on the $S_{2}$-particle surface increases weakly due to the increase in pressure gradient caused by viscoelasticity. Notably, the increasing presence of the Newtonian solvent in the dissolution, measured under $\beta$-increase, reduces the force intensity. For $De=1\times 10^{-1}$ and $\beta=0$, the viscoelastic lubrication force shows the largest rise relative to the Newtonian lubrication force, reaching an increase of up to $\approx 2\%$. One should bear in mind that in a steady state, the lubrication force exerted on the particles acts immediately. As discussed, the incorporation of viscoelasticity into the momentum balance equation via $N_{1}$, results in minor effects in the velocity, pressure-gradient, and pressure distributions. Thus, the effect of viscoelasticity on the lubrication force is small when the fluid response is considered instantaneous and when the effects of geometric changes derived from the particle approach are not considered.
Figure 4: Percent increase on the viscoelastic force with respect to a Newtonian response, against De-number under $\beta=\{0,0.2,0.5,0.8\}$.
@@ 4.1 Transient analysis
In the previous section, the steady-state case demonstrates the instantaneous response of viscoelastic fluids when subjected to a constant squeezing velocity. In this section, we introduce a time-dependent scenario, with the objective of examining the interplay between the particle dynamics and the time response of viscoelastic fluids characterised with the Oldroyd-B model. Therefore, we analyse the transient case by retaining the time-dependent terms in the stress components (11)-(14), which renders the following partial differential-equation system:
@@
For the corresponding initial conditions, we consider a completely relaxed stress, as follows:
@@
As particle $S_{2}$ moves with constant velocity $V$ towards particle $S_{1}$, its contour location changes with time, such that: $z_{2}(r,t)=h_{0}+\frac{r^{2}}{2a}-Vt$. Therefore, the dimensionless interparticle distance reduces with time as follows:
$H(R,T)=\frac{z_{2}(r,t)-z_{1}(r)}{h_{0}}=\frac{h(r,t)}{h_{0}}=1+\frac{R^{2}}{2\kappa_{a}}-T.$ (44)
@@ 4.2 Transient solutions
Substituting (46) in (40) and solving for the dimensionless shear stress under the initial condition in (43), yields:
@@
Differentiation with respect to $Z$ of (49), grouping of similar terms, and considering the following identities:
@@
the momentum balance equation in $r-$direction results in:
@@
The momentum equation in (51) relies on temporal functions $\left(\alpha(T),\frac{\partial P^{\prime}}{\partial R}(T)\right)$ that come from the viscoelastic stress. These terms emerge from the time-dependent development of the velocity field, as $S_{2}$ contour relocates over time. Integration of (51) under the boundary conditions in (9), yields the following velocity profile:
@@
Here, the temporal functions $\left(\alpha(T),\frac{\partial P^{\prime}}{\partial R}(T)\right)$ have the following limiting values:
@@
From the these limits in (53) and (54), one may conclude that:
@@
    which comes from (51) when viscoelastic effects are null, and represents the force balance for a Newtonian fluid of viscosity $\eta_{0}=\eta_{s}$. Therefore, the viscoelastic behaviour starts from a Newtonian fluid response.
@@
  • In the range $10<T<$ 1, the viscoelastic fluid response may evolve from a Newtonian fluid of viscosity $\eta_{0}=\eta_{s}$ to a viscoelastic fluid of viscosity $\eta_{0}=\eta_{s}+\eta_{p}$, whose force balance is given by (51). The dimensionless time expected to reach a steady value of $(g(T),f(T))$ is around $T\sim 0.5$ even for $De=1\times 10^{-1}$ (see Figure 5). However, since $Z_{2}(R,T)$ varies with time, the distance between the spheres will diminish and generate geometric changes (Ruangkriengsin et al., 2024), indicating that they are moving closer towards touching each other.
@@
  • In the limit of $T\rightarrow 1$, (54) simplifies to a version resembling the steady-state scenario in (30), that is, a viscoelastic fluid with a viscosity of $\eta_{0}=\eta_{s}+\eta_{p}$. Here, the contour of sphere $S_{2}$ continues to draw nearer to sphere $S_{1}$, which promotes the fluid to get confined.
@@
Figure 5: Temporal functions $g(T)$ and $f(T)$ originating from viscoelastic stress under $De=\{1\times 10^{-2},1\times 10^{-1}\}$ and $\beta=0.5$.
@@
Therefore, the response at relatively short times resembles the response of a Newtonian fluid with viscosity $\eta_{s}$, indicating that the parameter $\beta$ modulates the initial lubrication force. In the case of an upper-convected Maxwell fluid, i.e., for $\beta=0$, the force will start from zero, and, for $\beta=1$, the force will align with the expected response of a Newtonian fluid. The computations for pressure-gradient, pressure, and normal lubrication force follow the same numerical techniques outlined in Section 3.1, adding the functionality with time through $Z_{2}(R,T)$, $g(T)$ and $f(T)$. Considering the time in the domain $(0\leq T<1)$ in the transient case, the viscoelastic lubrication force is expected to increase over time or increase as distance is reduced, according with (44). In order to detect a significant alteration in distance, time needs to exceed $0.9$, such that $T>0.9$.
@@
We compare our solutions with expressions (52) and (54) in Zheng et al. (2023), corresponding to the pressure and the lubrication force, respectively. The main distinction between our formulation and that of Zheng et al. (2023), lies in their prescription of a Newtonian velocity profile, a stress-based boundary condition to compute the pressure-gradient, and a fluid characterised with the Maxwell model is considered. Figure 6 displays the time evolution of the pressure for different values of De-numbers comparing our solutions to those obtained by Zheng et al. (2023). For early times, specially under $T\lesssim 0.3$, our predictions yield pressures lower than those of the Newtonian reference solution. However, for $T\gtrsim 0.3$, the viscoelastic predictions converge to, and marginally exceed, the Newtonian result. In that sense, the development of the viscoelastic time-dependent functions in (51) promotes a rise both pressure and lubrication force. Noteworthy is the divergence of our solutions with respect to those of Zheng et al. (2023), which become apparent as $T\rightarrow 1$. Zheng et al. (2023) model predicts a decline in pressure, approaching negative values (see the inset in Figure 6), in contrast to the monotonic increase characteristic of the Newtonian solution, which is predicted with our model.
For the normal lubrication force in Figure 7, our solutions are in close agreement with those by Zheng et al. (2023) at short times. In addition, both models follow the Newtonian upper limit for all De-numbers. Nonetheless, for $T\gtrsim 0.2$, Zheng et al. (2023) model predicts a lubrication force that exceeds our solutions for equivalent De-levels.
@@
Figure 6: Dimensionless pressure against time under $De=\{1\times 10^{-6},5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},1\times 10^{-1}\}$ for $\epsilon=2.5\times 10^{-3}$. Symbols identify our numerical solutions whilst lines come from Eq.(52) in Zheng et al. (2023).
@@
Figure 7: Dimensionless lubrication force against time under $De=\{1\times 10^{-6},5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},1\times 10^{-1}\}$ for $\epsilon=2.5\times 10^{-3}$. Symbols identify our numerical solutions whilst lines come from Eq.(52) in Zheng et al. (2023).